OpenStudy (awkwardpanda):

In triangle DEF, m angle D= 43º, m angle E=62º, and EF=22 in. What is DE to the nearest tenth of an inch? a. 28.5 in b. 15.6 in c. 14.1 in d. 31.2 in

OpenStudy (juicstice):

set it up like a proportion

OpenStudy (awkwardpanda):

proportion?

OpenStudy (juicstice):

so D is the angle opposite of EF and F is the angle opposite of DE so first you have to find F

OpenStudy (juicstice):

180-105=75 so that is angle F to find DE you have to set up a proportion like this 43/22=75/x cross multiply

OpenStudy (juicstice):

22 x 75 =21650 then divide that by 43

OpenStudy (juicstice):

1650*

OpenStudy (awkwardpanda):

You're a lifesaver, thank you.

OpenStudy (juicstice):

you're welcome

OpenStudy (awkwardpanda):

I got 38. But that can't be right.

OpenStudy (juicstice):

I used the wrong thing I'm sorry. hold on I got it though

OpenStudy (awkwardpanda):

It's okay (:

OpenStudy (juicstice):

okay I got it a/sinA=b/sinB=c/sinC the capital letter is the angle the lower case is the side

OpenStudy (awkwardpanda):

sweet. how would i plug that in? i keep forgetting .___.

OpenStudy (juicstice):

okay so it would be like this d/sinD=f/sinF side DE=what we don't know so that would side EF=22 angle D=62 angle F=75 do you think you can plug that in

OpenStudy (awkwardpanda):

I'll try :3

OpenStudy (awkwardpanda):

d/sin(62 = f/sin(75)

OpenStudy (juicstice):

f=22

OpenStudy (juicstice):

d/sin(62)=22/sin(75)

OpenStudy (juicstice):

then just multiply 22 x sin(62) then divide that by sin(75)

OpenStudy (juicstice):

no wait f= is what we don't know so d=22

OpenStudy (juicstice):

sorry I'm just kinda tired

OpenStudy (juicstice):

22 x sin(75) divide that by sin(62)

OpenStudy (awkwardpanda):

same... i'm behind in math so i'm paying for it ha. i'll let you sleep. i'm sure i can figure it out, thank you.

OpenStudy (juicstice):

okay

OpenStudy (anonymous):

im gonna change DEF to ABC respectively just because i find it easier to visualize it that way... so basically you know m of angle A is 43 degrees, the m of angle B is 62 degrees, and the line from E to F (aka d) = 22 in. correct?

OpenStudy (anonymous):

in that case you would do another version of cross multiplication using sine. so you would put the sin of 43 or sin43/22 because those two are the A and the little a and make that multiplied with the sin of B or sin62/(x) so then you have the sin43= sin62*22 which is....

OpenStudy (anonymous):

some long number that i'm too tired to type out anyways so then you divide that number by the sin of 43 and you get about 28.5 by my calculations

OpenStudy (anonymous):

so A is your answer. hope that made sense i'm pretty tired XD

OpenStudy (anonymous):

the way i write it is sort of strange basically you have sin62/x=sin43/22. The easier way for me to look at it is as x/sin62=22/sin43 so then you just do 22sin62/sin43=x and simplify

OpenStudy (anonymous):

i hope that's a bit clearer

OpenStudy (awkwardpanda):

THANK YOU! :D

OpenStudy (anonymous):

sure thing :)

OpenStudy (anonymous):

i just finished this lesson haha so i remember how to do it XD

OpenStudy (anonymous):

any other ones you need help on?

OpenStudy (awkwardpanda):

yooo, so you did the unit test? YES I DO OML

OpenStudy (awkwardpanda):

Half of them are chinese to me xD

OpenStudy (anonymous):

oh boy... haha

OpenStudy (awkwardpanda):

wait wait, i might be exaggerating.

OpenStudy (anonymous):

*thinks about it* nah...im not exaggerating XD

OpenStudy (awkwardpanda):

number 8 (hopefully we have the same question) it says. find the height of the triangle. round the answer to one decimal place.

OpenStudy (awkwardpanda):

I swear i thought i had this, then bam i forgot .___.

OpenStudy (anonymous):

nope, dont have that. could you send a picture?

OpenStudy (awkwardpanda):

pickles. sure hold on.

OpenStudy (anonymous):

haha sure thing

OpenStudy (awkwardpanda):

Swear I'm not cheating but I'm so lost.

OpenStudy (anonymous):

mmk so same thing here. for the sake of understanding me im going to say the measure of angle a is 50 degrees, the measure of angle B is 90 degrees, the measure of angle C (which you dont actually need to know here) is 40 degrees, little a is x and little b is 10. you with me so far?

OpenStudy (anonymous):

from there, set it up like we did last time with big letter over little letter for what we know with all degrees being the big letters and all the segments being the little ones

OpenStudy (anonymous):

so you have sin50/x=sin90/10

OpenStudy (anonymous):

i hate looking at it like that so for me it's easier to see x/sin50=10/sin90 which is for our purposes the same exact thing. so you multiply both sides by sin50 to get 10sin50/sin90=x

OpenStudy (awkwardpanda):

wowzers, im writing this down

OpenStudy (anonymous):

if my calculator is right then you get about 7.7 and im not going to question my calculator... XD

OpenStudy (anonymous):

i hope you can see that XD my handwriting is kinda small

OpenStudy (awkwardpanda):

so is mine xD i can see it

OpenStudy (anonymous):

mmk so does that make sense?

OpenStudy (anonymous):

i could find little c the same way but im lazy and if im not asked to do it for credit then i wont haha

OpenStudy (awkwardpanda):

it makes sense, i wont ask you to find the c xD

OpenStudy (anonymous):

YAY!! :D

OpenStudy (anonymous):

ok so next one?

OpenStudy (anonymous):

well actually, why dont you just find the little c and ill do it too and that way you can make sure you get it.

OpenStudy (awkwardpanda):

i'll do it after this test, it's 1 am here xD i have a few more questions to finish. okay next question

OpenStudy (anonymous):

haha ok

OpenStudy (awkwardpanda):

for number 11, D is 68.8m^2

OpenStudy (anonymous):

its 1 here too but...insomnia...so ya doesnt really bother me much anymore XD

OpenStudy (anonymous):

your tests are awfully annoying arent they.... XD

OpenStudy (awkwardpanda):

Yes!! so very annoying ;-; i have a final coming up next so basically I'm about to die an annoying death.

OpenStudy (anonymous):

haha ill join ya.

OpenStudy (anonymous):

ok so i would first try to find little a or line segment bc because that makes life easier

OpenStudy (anonymous):

so a^2 + b^2 = c^2 so you have 23^2 = 7^2 + a^2

OpenStudy (anonymous):

= 529 = 49 + a^2. subtract the 49s and you have 480=a^2 rad of 480 is about 21.9

OpenStudy (anonymous):

with me so far?

OpenStudy (awkwardpanda):

yeah :D

OpenStudy (anonymous):

ok yay. so then you plug in what you know using the big letter over little letter formula (forgot the actual name XD) so you have sin(x)/21.9 = sin90/23

OpenStudy (anonymous):

so you have sin(x) = 21.9sin90/23

OpenStudy (anonymous):

(quick tip the sine of 90 is 1 so anything in front of a sin(90) is that number so in this case you just have 21.9/23)

OpenStudy (anonymous):

which is some long complicated number and im lazy XD ok so from there, you have to do the arc sin of that number or sin^(-1) to find your degree

OpenStudy (anonymous):

my calculator says 72.2

OpenStudy (anonymous):

but 72.3 also works

OpenStudy (anonymous):

make sense?

OpenStudy (awkwardpanda):

strangely enough yes

OpenStudy (anonymous):

haha yay

OpenStudy (awkwardpanda):

so i've been getting them right xD

OpenStudy (awkwardpanda):

number 11? i'm not so sure on that one.

OpenStudy (anonymous):

mmk

OpenStudy (anonymous):

fudge...this uses the longer equation...hold on...

OpenStudy (awkwardpanda):

I KNEW IT. Pickles

OpenStudy (anonymous):

a^2 = b^2 + c^2 - 2(b)(c)cosA is your equation and you can move those values around if your equation asks it

OpenStudy (anonymous):

your problems are not enjoyable at 1 in the morning my friend

OpenStudy (anonymous):

here we go

OpenStudy (awkwardpanda):

i feel.

OpenStudy (awkwardpanda):

wait, it's gotta be C. 48.2

OpenStudy (anonymous):

what is option D?

OpenStudy (awkwardpanda):

68.8m^2

OpenStudy (anonymous):

not so sure about that but interwebs....

OpenStudy (anonymous):

in which case A is your closest answer